Given f³ (x) = x2sin(x²), compute the derivative of f with respect to x.


1
Expert's answer
2020-04-30T19:36:43-0400

"f^{3}(x)=x^{2}\sin(x^{2})
\\
\implies
f(x)=(x^{2}\sin(x^{2}))^{\frac{1}{3}}
\\
\implies
f^{\prime}(x)=\frac{1}{3}(x^{2}\sin(x^{2}))^{\frac{-2}{3}}. \\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ . \ [2x \sin(x^{2})
+2x^{3} \cos(x^{2})]"